Read the following prompt and then answer the following questions:

29. You work as a financial analyst for a new subsidiary of Coca-Cola and you are analyzing the new firm’s capital structure. As you move through the data you try to answer the following to your boss

  1. You propose using no debt in the capital structure . Total invested capital is 50 million, Net income is 8 million and outstanding shares are 10 million. What are the EPS and ROE respectively?
    1. $.40 and 16%
    2. $1 and 20%
    3. $.80 and 16%
    4. cannot be determined
  2. Now assume there is 30% debt in the capital structure at a 5% coupon rate which makes net income 7.4 million, and this will reduce the outstanding shares and equity proportionately. What are the EPS and ROE now?
    1. $.80 and 16%
    2. b.$1.06 and 16%
    3. $.80 and 21%
    4. $1.06 and 21%
  3. Which statement is most true
    1. Adding debt makes ROE numbers lower but lowers the WACC
    2. Adding debt can increase EPS but may increase the WACC as more debt is added
    3. Adding debt can increase the ROE and will always result in the highest stock price and lowest WACC
  4. The new firm wants to minimize fixed costs. What does this mean?
    1. This will keep operating leverage down which in turn will lower business risk
    2. This will lower financial risk which in turn lowers the risk of the firm.
    3. Higher fixed costs will leverage up returns and lower business risk
    4. Higher operating leverage is directly associated with higher financial risk.
Answer #1

20.a)

EPS= 8/10=0.8 and ROE=8/50=16%

Option c is correct

b) 30% debt means $15mn Debt and $35 mn equity. Outstanding shares are 7million.

EPS=7.4/7=$1.06/share and ROE= 7.4/35=21%

Option d is correct

c) b.

  1. Adding debt can increase EPS but may increase the WACC as more debt is added

d.  

  1. This will keep operating leverage down which in turn will lower business risk